# A Novel State-Centric Necessary Condition for Time-Optimal Control of Controllable Linear Systems Based on Augmented Switching Laws

### v1

@@ V-C Necessary Condition for the Optimal Trajectory
Based on Theorem 2, this section provides an approach to constructing a feasible disturbed trajectory with the same augmented switching law. According to Theorem 1, this section shows that the existence of such a disturbed trajectory implies that the original trajectory is not optimal, resulting in a state-centric necessary for non-chattering optimal control of problem (1a).
@@
The Jacobian matrix of equality constraints induced by the augmented switching law can be calculated by Proposition 7.
@@
$\frac{\partial{\boldsymbol{x}}_{i}}{t_{j}}=\begin{dcases}\prod_{k=i}^{j+1}{\mathrm{e}}^{{\boldsymbol{A}}_{k}\left(t_{k}-t_{k-1}\right)}\left(\left({\boldsymbol{A}}_{j}-{\boldsymbol{A}}_{j+1}\right){\boldsymbol{x}}_{j}+{\boldsymbol{b}}_{j}-{\boldsymbol{b}}_{j+1}\right),\\ \qquad j<i,\\ {\boldsymbol{A}}_{i}{\boldsymbol{x}}_{i}+{\boldsymbol{b}}_{i},\,j=i,\\ 0,\,j>i,\end{dcases}$ (45)
where $\prod_{k=i}^{j+1}{\mathrm{e}}^{{\boldsymbol{A}}_{k}\left(t_{k}-t_{k-1}\right)}$ refers to
${\mathrm{e}}^{{\boldsymbol{A}}_{i}\left(t_{i}-t_{i-1}\right)}{\mathrm{e}}^{{\boldsymbol{A}}_{i-1}\left(t_{i-1}-t_{i-2}\right)}\dots\mathrm{e}^{{\boldsymbol{A}}_{j+1}\left(t_{j+1}-t_{j}\right)}.$ (46)
@@
Evidently, $\forall j>i$, ${\boldsymbol{x}}_{i}$ is independent with $t_{j}$; hence, $\frac{\partial{\boldsymbol{x}}_{i}}{t_{j}}={\boldsymbol{0}}$. According to Lemma 1,
${\boldsymbol{x}}_{i}={\mathrm{e}}^{{\boldsymbol{A}}_{i}\left(t_{i}-t_{i-1}\right)}{\boldsymbol{x}}_{i-1}+\int_{t_{i-1}}^{t_{i}}{\mathrm{e}}^{{\boldsymbol{A}}_{i}\left(t_{i}-\tau\right)}{\boldsymbol{b}}_{i}\mathrm{d}\tau.$ (47)
Therefore, $\forall i\in\left[M\right]$,
$\frac{\partial{\boldsymbol{x}}_{i}}{\partial t_{i}}={\boldsymbol{A}}_{i}{\boldsymbol{x}}_{i}+{\boldsymbol{b}}_{i}.$ (48)
Applying (48), $\forall i>1$, it has
@@
$\forall i>1$, $j<i-1$, it has
$\frac{\partial{\boldsymbol{x}}_{i}}{\partial t_{j}}=\prod_{k=i}^{j+2}{\mathrm{e}}^{{\boldsymbol{A}}_{k}\left(t_{k}-t_{k-1}\right)}\frac{\partial{\boldsymbol{x}}_{j+1}}{\partial t_{j}}$ (50)
$=$ $\prod_{k=i}^{j+1}{\mathrm{e}}^{{\boldsymbol{A}}_{k}\left(t_{k}-t_{k-1}\right)}\left(\left({\boldsymbol{A}}_{j}-{\boldsymbol{A}}_{j+1}\right){\boldsymbol{x}}_{j}+{\boldsymbol{b}}_{j}-{\boldsymbol{b}}_{j+1}\right).$
Therefore, (45) holds. ∎
Utilizing the notations in Section V-B, the augmented switching law provides a system of equalities on every keypoints. Denote that $\forall i\in\left[M\right]$, ${\boldsymbol{F}}_{i}{\boldsymbol{x}}_{i}+{\boldsymbol{g}}_{i}={\boldsymbol{0}}$ and ${\boldsymbol{C}}_{i}{\boldsymbol{x}}_{i}+{\boldsymbol{d}}_{i}<{\boldsymbol{0}}$, where ${\boldsymbol{x}}_{i}={\boldsymbol{x}}\left(t_{i}\right)$ and ${\boldsymbol{F}}_{i}$ has full row rank. For a constrained arc with active constraints $\mathcal{P}$, the equality constraints induced by $\mathcal{P}$ are added in only one end instead of both ends of the arcs since one end satisfies $\mathcal{P}$ implies that the whole arc satisfies $\mathcal{P}$. Note that the additional end-constraints at both ends are eliminated by equality constraints induced by $\mathcal{P}$. Similar processes are applied if $u\equiv{u_{\mathrm{m}}}$ or $u\equiv-{u_{\mathrm{m}}}$ during a constrained arc. ${\boldsymbol{F}}_{i}{\boldsymbol{x}}_{i}\left({\boldsymbol{t}}\right)+{\boldsymbol{g}}_{i}$ for $i\in\left[M\right]$ induce a function ${\boldsymbol{H}}\left({\boldsymbol{t}}\right)$. Then, Theorem 3 provides a necessary condition for the optimal trajectory.
Before providing Theorem 3, the implicit function theorem [29] should be introduced.
Lemma 3 (The Implicit Function Theorem [29]).
Assume $S\subset{\mathbb{R}}^{p+q}$ is open and non-empty. ${\boldsymbol{F}}:S\to{\mathbb{R}}^{q}$ is $C^{1}$ continuous. $\left({\boldsymbol{\xi}}_{0},{\boldsymbol{\eta}}_{0}\right)\in S$, satisfying ${\boldsymbol{F}}\left({\boldsymbol{\xi}}_{0},{\boldsymbol{\eta}}_{0}\right)={\boldsymbol{0}}$ and $\det\frac{\partial{\boldsymbol{F}}}{\partial{\boldsymbol{\eta}}}\left({\boldsymbol{\xi}}_{0},{\boldsymbol{\eta}}_{0}\right)\not=0$. Then, $\exists\delta_{1},\delta_{2}>0$, s.t. $\forall{\boldsymbol{\xi}}\in B_{\delta_{1}}\left({\boldsymbol{\xi}}_{0}\right)$, $\exists!{\boldsymbol{\eta}}\in B_{\delta_{2}}\left({\boldsymbol{\eta}}_{0}\right)$ satisfying ${\boldsymbol{F}}\left({\boldsymbol{\xi}},{\boldsymbol{\eta}}\right)={\boldsymbol{0}}$. The above relation induces a mapping ${\boldsymbol{f}}:B_{\delta_{1}}\left({\boldsymbol{\xi}}_{0}\right)\to B_{\delta_{2}}\left({\boldsymbol{\eta}}_{0}\right)$. Then, ${\boldsymbol{f}}$ is $C^{1}$ continuous, where $\frac{\mathrm{d}{\boldsymbol{f}}}{\mathrm{d}{\boldsymbol{\xi}}}\left({\boldsymbol{\xi}}\right)=-\frac{\partial{\boldsymbol{F}}}{\partial{\boldsymbol{\eta}}}\left({\boldsymbol{\xi}},{\boldsymbol{f}}\left({\boldsymbol{\xi}}\right)\right)^{-1}\frac{\partial{\boldsymbol{F}}}{\partial{\boldsymbol{\xi}}}\left({\boldsymbol{\xi}},{\boldsymbol{f}}\left({\boldsymbol{\xi}}\right)\right)$.
@@
Assume that the optimal trajectory ${\boldsymbol{x}}={\boldsymbol{x}}^{*}\left(t\right)$ of problem (1a) satisfies Assumption 1. Denote the equality constraints induced by the augmented switching law as ${\boldsymbol{H}}\left({\boldsymbol{t}}^{*}\right)={\boldsymbol{0}}$, where ${\boldsymbol{t}}^{*}=\left(t_{i}^{*}\right)_{i=1}^{M}$ is the arriving time of each keypoints and $t_{0}=0$. Then, $\frac{\partial{\boldsymbol{H}}}{\partial{\boldsymbol{t}}_{1:\left(M-1\right)}}\left({\boldsymbol{t}}^{*}\right)$ does not have full row rank.
@@
Assume that the $i$-th row of $\frac{\partial{\boldsymbol{H}}}{\partial{\boldsymbol{t}}_{1:\left(M-1\right)}}\left({\boldsymbol{t}}^{*}\right)$ is linearly independent where $i\in\mathcal{I}\subset\left[M-1\right]$ and #$\#\mathcal{I}=M^{\prime}$. Let $\varepsilon>0$ be small enough to satisfy the condition of Theorem 2 and Lemma 3. The implicit function ${\boldsymbol{f}}:\left(t_{i}\right)_{i\in{\left.\left[M\right]\middle\backslash\mathcal{I}\right.}}\mapsto\left(t_{i}\right)_{i\in\mathcal{I}}$ is induced by ${\boldsymbol{H}}\left({\boldsymbol{t}}\right)={\boldsymbol{0}}$. Based on ${\boldsymbol{f}}$, $\exists{\boldsymbol{t}}^{\prime}$ satisfies ${\boldsymbol{H}}\left({\boldsymbol{t}}^{\prime}\right)={\boldsymbol{0}}$, $t_{M}^{\prime}<t_{M}^{*}$, and $\left\|{\boldsymbol{t}}^{\prime}-{\boldsymbol{t}}^{*}\right\|<\varepsilon$. According to Theorem 2, the disturbed trajectory ${\boldsymbol{x}}={\boldsymbol{x}}^{\prime}\left(t\right)$ is feasible since ${\boldsymbol{H}}\left({\boldsymbol{t}}^{\prime}\right)={\boldsymbol{0}}$ implies that ${\boldsymbol{x}}^{\prime}$ and ${\boldsymbol{x}}^{*}$ have the same augmented switching law. However, ${\boldsymbol{x}}^{\prime}$ achieves a shorter terminal time $t_{M}^{\prime}<t_{M}^{*}$, which contradicts the optimality of ${\boldsymbol{x}}^{*}$. ∎
@@
In fact, if a given non-chattering Bang-Singular-Bang trajectory does not satisfy the necessary condition in Theorem 3, then one can perform numerical local optimization of trajectories based on Theorem 3. For example, let $t_{M}$ be the optimization objective and consider the equality and inequality constraints of the augmented switching law as the constraints of the optimization problem. In this way, it is hopeful to obtain a locally optimal solution or a stationary point of problem (1a) through iteratively solving an $M$-dimensional nonlinear problem. Note that when applying shooting methods [30], problem (1a) is highly nonlinear and non-convex, and the long motion time may result in unacceptable computation cost. In contrast, the local optimization method on ${\boldsymbol{t}}$ based on Theorem 1a will be more efficient due to the significantly lower dimensions and the suboptimal feasible initial value. The above method will be developed in future works.
Theorem 3 provides a necessary condition for the optimal control of problem (1a) under Assumption 1. Different from the existing costate-based necessary conditions [14], Theorem 3 requires no information on the costates, which significantly simplifies the computation of the necessary conditions. Compared to existing state-centric necessary conditions [22], Theorem 3 is effective in theoretical reasoning and has a lower computational complexity in numerical computation.
Some further discussions on the Jacobian matrix are necessary. Consider a keypoint $t_{i}$ with a constraint ${\boldsymbol{c}}^{\top}{\boldsymbol{x}}+d\leq{\boldsymbol{0}}$ active at $t_{i}$. Assume that the constraint is of $\hat{r}$-th order w.r.t the arc $\left(t_{i-1},t_{i}\right)$ where $\dot{\boldsymbol{x}}={\boldsymbol{A}}_{i}{\boldsymbol{x}}+{\boldsymbol{b}}_{i}$. In other words, if $1\leq\hat{r}<n$, then $\left(-1\right)^{\hat{r}}\frac{\mathrm{d}^{\hat{r}}\left({\boldsymbol{c}}^{\top}{\boldsymbol{x}}+d\right)}{\mathrm{d}t^{\hat{r}}}\left(t_{i}^{-}\right)>0$ and $\forall 1\leq r<\hat{r}$, $\frac{\mathrm{d}^{r}\left({\boldsymbol{c}}^{\top}{\boldsymbol{x}}+d\right)}{\mathrm{d}t^{r}}\left(t_{i}^{-}\right)=0$. If $\hat{r}=n$, then $\forall r\in{\mathbb{N}}^{*}$, $\frac{\mathrm{d}^{r}\left({\boldsymbol{c}}^{\top}{\boldsymbol{x}}+d\right)}{\mathrm{d}t^{r}}\left(t_{i}^{-}\right)=0$. Among them. $\forall r\in{\mathbb{N}}^{*}$, $\frac{\mathrm{d}^{r}\left({\boldsymbol{c}}^{\top}{\boldsymbol{x}}+d\right)}{\mathrm{d}t^{r}}\left(t_{i}^{-}\right)={\boldsymbol{c}}^{\top}{\boldsymbol{A}}_{i}^{r-1}\left({\boldsymbol{A}}_{i}{\boldsymbol{x}}_{i}+{\boldsymbol{b}}_{i}\right)$. According to Proposition 7, $\forall r\in{\mathbb{N}}$,
$\frac{\partial}{\partial t_{i}}\left.\frac{\mathrm{d}^{r}\left({\boldsymbol{c}}^{\top}{\boldsymbol{x}}+d\right)}{\mathrm{d}t^{r}}\right|_{t_{i}^{-}}=\frac{\partial}{\partial t_{i}}\left({\boldsymbol{c}}^{\top}{\boldsymbol{A}}_{i}^{r-1}\left({\boldsymbol{A}}_{i}{\boldsymbol{x}}_{i}+{\boldsymbol{b}}_{i}\right)\right)$ (52)
$=$ $\left({\boldsymbol{c}}^{\top}{\boldsymbol{A}}_{i}^{r}\left({\boldsymbol{A}}_{i}{\boldsymbol{x}}_{i}+{\boldsymbol{b}}_{i}\right)\right)=\left.\frac{\mathrm{d}^{r+1}\left({\boldsymbol{c}}^{\top}{\boldsymbol{x}}+d\right)}{\mathrm{d}t^{r+1}}\right|_{t_{i}^{-}}.$
In other words, the Jacobian matrix of ${\boldsymbol{c}}^{\top}{\boldsymbol{x}}+d\leq{\boldsymbol{0}}$ w.r.t. $t_{i}$ is $\left.\frac{\mathrm{d}^{r+1}\left({\boldsymbol{c}}^{\top}{\boldsymbol{x}}+d\right)}{\mathrm{d}t^{r+1}}\right|_{t_{i}^{-}}{\boldsymbol{e}}_{\hat{r}}$. Specifically, if $\hat{r}=n$, then the Jacobian matrix of ${\boldsymbol{c}}^{\top}{\boldsymbol{x}}+d\leq{\boldsymbol{0}}$ w.r.t. $t_{i}$ is ${\boldsymbol{0}}$.
VI Applications in Chain-of-Integrators Systems and Simulation Experiments
This section provides some applications of the state-centric necessary condition, i.e., Theorem 3, for the optimal control of problem (1a). For convenience, the chain-of-integrators system with box constraints is considered in this section, where the notations have been introduced in Section IV-C.
Time-optimal control for chain-of-integrators systems with box constraints is an open and challenging problem in the optimal control domain, yet to be resolved. The problem can be summarized as follows:
@@ VI Applications in Chain-of-Integrators Systems and Simulation Experiments
${\mathrm{s.t.}}\quad$ $\dot{x}_{k}\left(t\right)=x_{k-1}\left(t\right),\,\forall 1<k\leq n,\,t\in\left[0,t_{\mathrm{f}}\right],$ (53b)
@@
$\left|x_{k}\left(t\right)\right|\leq x_{{\mathrm{m}}k},\,\forall 1\leq k\leq n,\,t\in\left[0,t_{\mathrm{f}}\right],$ (53e)
@@
The Jacobian matrix, except the last column induced by the augmented switching law, denoted as ${\boldsymbol{J}}$, can be calculated as follows:
Corollary 1.
Assume that $\forall i\in\left[N\right]$, $u\equiv u_{i}\in\left\{\pm{u_{\mathrm{m}}},0\right\}$ in $\left(t_{i-1},t_{i}\right)$, and denote ${\boldsymbol{x}}_{i}={\boldsymbol{x}}\left(t_{i}\right)$. Then,
$\frac{\partial{\boldsymbol{x}}_{i}}{\partial t_{j}}=\begin{dcases}-\Delta u_{j+1}{\boldsymbol{\phi}}\left(t_{i}-t_{j}\right),&j<i\\ \left[\begin{array}[]{c}u_{i}\\ {\boldsymbol{x}}_{i,1:\left(n-1\right)}\end{array}\right],&j=i\\ {\boldsymbol{0}},&j>i.\end{dcases}$ (54)
Among them, $\Delta u_{i}=u_{i}-u_{i-1}$ and $u_{0}=0$. ${\boldsymbol{\phi}}\left(t\right)\triangleq\left(\phi_{k}\left(t\right)\right)_{k=1}^{n}$ where $\phi_{k}\left(t\right)=\frac{t^{k-1}}{\left(k-1\right)!}$.
Proof.
Proposition 7 directly implies Corollary 1. ∎

### v2

@@ V-C State-Centric Necessary Condition of Optimality
Based on Theorem 2, this section provides an approach to constructing a feasible perturbed trajectory with the same ASL. According to Theorem 1, this section shows that the existence of such a perturbed trajectory implies that the original trajectory is not optimal, resulting in a state-centric necessary for non-chattering optimal control of problem (1a).
@@
The Jacobian matrix of equality constraints induced by the ASL can be calculated through Proposition 8.
@@
$\frac{\partial{\boldsymbol{x}}_{i}}{\partial t_{j}}=\begin{dcases}{\boldsymbol{P}}_{j,i}\left(\left({\boldsymbol{A}}_{j}-{\boldsymbol{A}}_{j+1}\right){\boldsymbol{x}}_{j}+{\boldsymbol{b}}_{j}-{\boldsymbol{b}}_{j+1}\right),\,j<i,\\ {\boldsymbol{A}}_{i}{\boldsymbol{x}}_{i}+{\boldsymbol{b}}_{i},\,j=i,\\ {\boldsymbol{0}},\,j>i,\end{dcases}$ (45)
where ${\boldsymbol{P}}_{j,i}={\mathrm{e}}^{{\boldsymbol{A}}_{i}\left(t_{i}-t_{i-1}\right)}{\mathrm{e}}^{{\boldsymbol{A}}_{i-1}\left(t_{i-1}-t_{i-2}\right)}\dots\mathrm{e}^{{\boldsymbol{A}}_{j+1}\left(t_{j+1}-t_{j}\right)}$.
@@
$\forall j>i$, $\frac{\partial{\boldsymbol{x}}_{i}}{\partial t_{j}}={\boldsymbol{0}}$ holds evidently. According to Lemma 1, ${\boldsymbol{x}}_{i}={\mathrm{e}}^{{\boldsymbol{A}}_{i}\left(t_{i}-t_{i-1}\right)}{\boldsymbol{x}}_{i-1}+\int_{t_{i-1}}^{t_{i}}{\mathrm{e}}^{{\boldsymbol{A}}_{i}\left(t_{i}-\tau\right)}{\boldsymbol{b}}_{i}\mathrm{d}\tau$; hence, $\forall i\in\left[M\right]$, $\frac{\partial{\boldsymbol{x}}_{i}}{\partial t_{i}}={\boldsymbol{A}}_{i}{\boldsymbol{x}}_{i}+{\boldsymbol{b}}_{i}$ holds. Then, $\forall i>1$, it holds that
@@
$\forall i>1$, $j<i-1$, $\frac{\partial{\boldsymbol{x}}_{i}}{\partial t_{j}}=\prod_{k=i}^{j+2}{\mathrm{e}}^{{\boldsymbol{A}}_{k}\left(t_{k}-t_{k-1}\right)}\frac{\partial{\boldsymbol{x}}_{j+1}}{\partial t_{j}}$ holds, which implies (45). ∎
Utilizing the notations in Section V-B, the ASL provides a system of equalities on every keypoints. Denote that $\forall i\in\left[M\right]$, ${\boldsymbol{F}}_{i}{\boldsymbol{x}}_{i}+{\boldsymbol{g}}_{i}={\boldsymbol{0}}$ and ${\boldsymbol{C}}_{i}{\boldsymbol{x}}_{i}+{\boldsymbol{d}}_{i}<{\boldsymbol{0}}$, where ${\boldsymbol{x}}_{i}={\boldsymbol{x}}\left(t_{i}\right)$ and ${\boldsymbol{F}}_{i}$ has full row rank. For a constrained arc with active constraints $\mathcal{P}$, the equality constraints induced by $\mathcal{P}$ are added in only one end instead of both ends of the arcs since one end satisfies $\mathcal{P}$ implies that the whole arc satisfies $\mathcal{P}$. Note that the additional end-constraints at both ends are eliminated by equality constraints induced by $\mathcal{P}$. Similar processes are applied if $u\equiv{u_{\mathrm{m}}}$ or $u\equiv-{u_{\mathrm{m}}}$ during a constrained arc. ${\boldsymbol{F}}_{i}{\boldsymbol{x}}_{i}\left({\boldsymbol{t}}\right)+{\boldsymbol{g}}_{i}$ for $i\in\left[M\right]$ induce a function ${\boldsymbol{H}}\left({\boldsymbol{t}}\right)$. Then, Theorem 3 provides a necessary condition of optimality.
@@
Assume that the optimal trajectory ${\boldsymbol{x}}={\boldsymbol{x}}^{*}\left(t\right)$ of problem (1a) satisfies Assumption 1. Denote the equality constraints induced by the ASL as ${\boldsymbol{H}}\left({\boldsymbol{t}}^{*}\right)={\boldsymbol{0}}$, where ${\boldsymbol{t}}^{*}=\left(t_{i}^{*}\right)_{i=1}^{M}$ is the arriving time of each keypoints and $t_{0}=0$. Then, $\frac{\partial{\boldsymbol{H}}}{\partial{\boldsymbol{t}}_{1:\left(M-1\right)}}\left({\boldsymbol{t}}^{*}\right)$ does not have full row rank.
@@
Assume that the $i$-th row of $\frac{\partial{\boldsymbol{H}}}{\partial{\boldsymbol{t}}_{1:\left(M-1\right)}}\left({\boldsymbol{t}}^{*}\right)$ is linearly independent where $i\in\mathcal{I}\subset\left[M-1\right]$ and $\left|\mathcal{I}\right|=M^{\prime}$. Let $\varepsilon>0$ be sufficiently small to satisfy the condition of Theorem 2 and the implicit function theorem [31]. The implicit function ${\boldsymbol{f}}:\left(t_{i}\right)_{i\in{\left.\left[M\right]\middle\backslash\mathcal{I}\right.}}\mapsto\left(t_{i}\right)_{i\in\mathcal{I}}$ is induced by ${\boldsymbol{H}}\left({\boldsymbol{t}}\right)={\boldsymbol{0}}$. Based on ${\boldsymbol{f}}$, $\exists{\boldsymbol{t}}^{\prime}$ satisfies ${\boldsymbol{H}}\left({\boldsymbol{t}}^{\prime}\right)={\boldsymbol{0}}$, $t_{M}^{\prime}<t_{M}^{*}$, and $\left\|{\boldsymbol{t}}^{\prime}-{\boldsymbol{t}}^{*}\right\|<\varepsilon$. According to Theorem 2, the perturbed trajectory ${\boldsymbol{x}}={\boldsymbol{x}}^{\prime}\left(t\right)$ is feasible since ${\boldsymbol{H}}\left({\boldsymbol{t}}^{\prime}\right)={\boldsymbol{0}}$ implies that ${\boldsymbol{x}}^{\prime}$ and ${\boldsymbol{x}}^{*}$ have the same ASL. However, ${\boldsymbol{x}}^{\prime}$ achieves a shorter terminal time $t_{M}^{\prime}<t_{M}^{*}$, which contradicts the optimality of ${\boldsymbol{x}}^{*}$. ∎
Remark.
In contrast to existing costate-based necessary conditions [9], Theorem 3 requires no information on costates, which significantly simplifies the computation. Compared to existing state-centric necessary conditions [24], Theorem 3 is effective in theoretical reasoning and has a lower computational complexity in numerical computation. Some examples are provided in Sections VI-B and VI-C.
@@
If a given non-chattering BBS trajectory does not satisfy the necessary condition in Theorem 3, then one can perform numerical optimization of the trajectory based on Theorem 3. Numerical examples are provided in Section VII.
VI Applications in Chain-of-Integrator Systems
This section provides some applications of the state-centric necessary condition, i.e., Theorem 3, for the optimal control of problem (1a). For convenience, the chain-of-integrator system with box constraints is considered in this section, where the notations have been introduced in Section IV-C.
Time-optimal control for chain-of-integrator systems with box constraints is an open and challenging problem in the optimal control domain, yet to be resolved. The problem can be summarized as follows:
@@ VI Applications in Chain-of-Integrator Systems
[…]
@@
[…]